OpenStudy (anonymous):

f '(x) of f(x)=8e^x-e^2x

OpenStudy (anonymous):

You are finding inverse of F(x) right? ( I just want to make sure)

OpenStudy (anonymous):

derivative

OpenStudy (anonymous):

iis =8e^x+2e^x2

OpenStudy (anonymous):

I need to get the derivative of f(x :)

OpenStudy (anonymous):

=8e^x - 2e^2x

OpenStudy (anonymous):

\[\frac{d}{dx}[e^u]=e^u(u')\]

OpenStudy (anonymous):

how chlorophyll did his

OpenStudy (anonymous):

you can pul the coefficients out if they bother you also so like \[8\frac{d}{dx}e^x\]